# CEFR-Annotated WordNet: LLM-Based Proficiency-Guided Semantic Database for Language Learning

### v1

@@ Abstract
Although WordNet is a valuable resource owing to its structured semantic networks and extensive vocabulary, its fine-grained sense distinctions can be challenging for second-language learners.
To address this, we developed a WordNet annotated with the Common European Framework of Reference for Languages (CEFR), integrating its semantic networks with language-proficiency levels.
@@
To validate our method, we constructed a large-scale corpus containing both sense and CEFR-level information from our annotated WordNet and used it to develop contextual lexical classifiers.
Our experiments demonstrate that models fine-tuned on our corpus perform comparably to those trained on gold-standard annotations.
Furthermore, by combining our corpus with the gold-standard data, we developed a practical classifier that achieves a Macro-F1 score of 0.81, indicating the high accuracy of our annotations.
Our annotated WordNet, corpus, and classifiers are publicly available to help bridge the gap between natural language processing and language education, thereby facilitating more effective and efficient language learning.
@@
| Masato Kikuchi✉††thanks: ✉ Corresponding author., Masatsugu Ono, Toshioki Soga˘, |
@@
| Nagoya Institute of Technology, Muroran Institute of Technology, |
| ˘Chitose Institute of Science and Technology, Hokkaido University |
| {kikuchi, ozono}@nitech.ac.jp, onomasa@muroran-it.ac.jp, |
@@ 1. Introduction
WordNet Fellbaum (1998) is a large-scale English lexical database that organizes approximately 155,000 words and 207,000 senses for nouns, verbs, adjectives, and adverbs into hierarchical semantic networks. It groups semantically similar words and links senses through rich relations such as hypernymy, hyponymy, synonymy, and antonymy. Because WordNet and its construction software are publicly available, they can be easily integrated into AI applications. Consequently, WordNet underpins a broad range of natural language processing (NLP) technologies—including machine translation Moussallem et al. (2018), semantic analysis Moskvoretskii et al. (2024), and natural language generation Vicente et al. (2014)—owing to its accessible interface and well-structured networks. These technologies also support computer-assisted language learning (CALL) by aiding in vocabulary acquisition, reading comprehension, writing assistance, automated question generation, and automated assessment.
@@
Although leveraging semantic networks can enhance foreign-language learning Kiritani et al. (2012), WordNet was not designed for educational purposes and presents challenges for second-language (L2) learners. A key issue is its overly fine-grained sense distinctions and the sheer number of senses for many words. This requires learners to identify the appropriate sense for a given context and proficiency level, which increases their cognitive load. While this problem is widely discussed in NLP-related literature Navigli (2006)Lacerra et al. (2020), it has received limited attention in language education. Our goal was to develop a novel version of WordNet and leverage the resulting technologies and resources to enhance language-learning efficiency. The first step involves adapting WordNet for L2 learners, bridging the gap between NLP lexical resources and language education. A previous work Kikuchi et al. (2024) clustered fine-grained WordNet sense definitions (glosses) from learner-oriented dictionaries. This study adopts a different approach by integrating Common European Framework of Reference for Languages (CEFR) proficiency levels into WordNet, enabling the presentation of senses that align with a learner’s proficiency. To build large-scale, practical resources, we employ a simple, large language model (LLM)-based method for efficient and accurate semantic annotation that reduces the time, labor, and cost associated with manual annotation. This automatic approach also ensures that our WordNet can be scaled flexibly.
@@
The CEFR is an international standard for describing language proficiency across six levels (A1, A2, B1, B2, C1, C2), from basic to advanced. Each level is defined by “can-do descriptors” that specify expected communication abilities. We used an LLM to annotate WordNet senses with these CEFR levels, thereby constructing a CEFR-annotated WordNet. As shown in Figure 1(a), these annotations can be used in conjunction with semantic networks to help learners acquire vocabulary while considering relationships among words, as well as to learn basic and advanced paraphrases through synonyms. The annotation pipeline involves three steps. First, we gather glosses for target words from WordNet and the English Vocabulary Profile (EVP) Online111https://englishprofile.org/?menu=evp-online Capel (2012), which provides CEFR levels for individual senses. Next, an LLM computes the semantic similarity between the glosses from WordNet and the EVP. Finally, we assign CEFR levels to the corresponding WordNet senses based on these similarity scores.
@@
As our method for annotating WordNet senses with CEFR levels is automatic, it eliminates the need for labor-intensive manual work. However, because automatic labels can contain errors and WordNet lacks gold-standard CEFR levels, their reliability must be verified indirectly. To address this, we built contextual CEFR-level classifiers that predict a sense’s proficiency level from its usage, as shown in Figure 1(b). These classifiers predict the CEFR level for a word sense based on its context, not just for the word itself. We assess the quality of our annotations by comparing classifiers trained on our data with those trained on the EVP gold-standard levels. We also examine the effectiveness of the LLMs for this task in few-shot and fine-tuning settings.
@@
CEFR-Annotated WordNet. We developed a new resource by assigning CEFR proficiency levels to 10,644 WordNet senses corresponding to 5,645 lemmas, effectively linking the WordNet sense inventory with CEFR standards. Our annotated WordNet covers approximately 80% of all single-word senses in the EVP (8,289 out of 10,394).
Prompt-Only LLM Annotation. We introduce a novel, automated method that leverages the semantic understanding of LLMs to assign CEFR levels to word senses. This is achieved by measuring the semantic similarity between WordNet glosses and EVP entries. The method, implemented entirely through prompting, is simple, reproducible, and significantly less costly than manual annotation. We also indirectly demonstrate that manual annotation tasks based on semantic matching can be automated with high accuracy.
SemCor-CEFR Corpus. Using our annotated WordNet, we assigned CEFR levels to the word senses in SemCor 3.0222http://lcl.uniroma1.it/wsdeval/training-data Miller et al. (1993), a widely used sense-annotated corpus. This resulted in a large-scale corpus with over 110,000 sense and level annotations across more than 5,500 WordNet senses. As modern NLP relies on large corpora for advanced training and analysis, our resource is a valuable contribution to NLP and educational-technology research.
Contextual Lexical CEFR-Level Classifiers. We demonstrate the validity of our CEFR-level annotations by training a classifier on our corpus that performs comparably to that trained on gold-standard EVP data. Additionally, by fine-tuning the LLM on both our annotated data and the gold-standard levels, we developed a practical classifier that achieves a Macro-F1 score of 0.81. Our analysis suggests that these classifiers can accurately predict CEFR levels in a broad range of contexts.
@@ 2.1. WordNets for Language Learning
As noted in the introduction, WordNet was not originally designed for educational use. To address this limitation, several learning-oriented WordNets have been developed for multiple languages Bosch and Griesel (2018), and numerous studies have explored their application in language learning Gonzalez-Dios (2019). Some of this work has focused on visualizing word hierarchies and semantic relations to aid learners Sun et al. (2011); Kiritani et al. (2012); Gawde et al. (2024). Others have tailored vocabulary, glosses, and usage examples to match learners’ proficiency levels Redkar et al. (2018); Osenova and Simov (2024). However, most of these efforts rely on manually curated resources and pay limited attention to word-sense information. By contrast, we introduce a novel approach that automatically annotates WordNet senses with proficiency levels. Our method can be integrated with existing techniques—such as semantic network visualization and multimodal WordNets Marciniak (2020)—to further enhance its utility in language-learning contexts.
@@ 2.2. Lexical Complexity Prediction
Lexical complexity prediction (LCP) North et al. (2023); Shardlow et al. (2024) has recently gained significant attention for estimating word complexity from context. In this field, “complexity”—which aligns with the CEFR levels in our study—is typically predicted as either binary (e.g., simple/complex) or on a continuous scale. Our work is closely related to SemEval-2021 Task 1 Shardlow et al. (2021), which shares a similar classification setting. For this task, the organizers released the CompLex 2.0 dataset333https://github.com/MMU-TDMLab/CompLex Shardlow et al. (2022), wherein words in context were rated by multiple annotators on a five-point Likert scale. The final scores are represented as a continuous value in , computed as the mean of these ratings. These continuous values can capture finer, context-driven differences compared with ordinal labels.
However, our approach differs from that of LCP in several key ways. First, LCP is primarily designed as a precursor to lexical simplification Paetzold and Specia (2017)—replacing complex words with simpler ones—rather than for explicitly presenting complexity information to L2 learners. Second, annotators for CompLex 2.0 were not provided with glosses; therefore, identical senses could receive different scores in different contexts. Third, the dataset is limited to 9,000 nouns, excluding other parts of speech (PoS).
By contrast, our method assigns a CEFR level to each word sense, adhering to an international proficiency standard. We extend this annotation to over 110,000 instances of nouns, verbs, adjectives, and adverbs in the large-scale SemCor corpus, making our resource more than ten times larger than CompLex 2.0. In Section 6.1, we analyze the correlation between the CompLex 2.0 complexity scores and the CEFR levels predicted by our models.
@@ 2.3. CEFR-Based Educational Technology
The CEFR is a foundational standard in educational technology, widely applied in the automatic assessment of short sentences Tack et al. (2017); Uchida et al. (2024), teaching materials Pilán et al. (2016), writing skills Kerz et al. (2021); Schmalz and Brutti (2021), and learner proficiency Gaillat et al. (2022). Recent research has also focused on the interaction between LLMs and the CEFR, exploring how well these models understand proficiency levels Benedetto et al. (2025) and how to control the difficulty of the vocabulary and sentences they generate Alfter (2024); Malik et al. (2024); Barayan et al. (2025). These efforts, along with the development of numerous CEFR-aligned lexical datasets, underscore the CEFR’s central role in the field.
For example, the CEFRLex project provides machine-readable lexical resources with word and sense frequency counts by CEFR level Pintard and François (2020) for English and other languages Dürlich and François (2018); François et al. (2014); Tack et al. (2018); François et al. (2016); Volodina et al. (2016). However, it does not assign a unique CEFR level to each sense, making it unsuitable for tasks requiring sense-specific proficiency annotations. The Sense Complexity Dataset (SeCoDa) Strohmaier et al. (2020) does provide sense-in-context CEFR annotations, but its sense labels are not aligned with those of WordNet, and its small size (1,432 words) limits its applicability within WordNet’s semantic framework. Our work addresses these gaps by annotating over 110,000 word instances with sense-specific CEFR levels, thereby substantially expanding the available data on lexical difficulty.
Despite progress in LCP, few studies have focused on classifying vocabulary into CEFR levels based on context. Aleksandrova and Pouliot (2023) proposed ME6 Contextual, a bidirectional encoder representations from transformers (BERT) Devlin et al. (2019)-based classifier that, similar to our models, is trained on a CEFR-annotated corpus to directly predict a word’s level from its context. This direct-prediction approach allows it to classify words not seen during training. By contrast, Bannò et al. (2025) introduced an indirect-prediction method wherein an LLM selects the appropriate EVP sense for a word in context before mapping it to a CEFR level. The performance of this approach depends on the quality of cues from the data source. To isolate the effect of different data sources on classification performance, we reimplemented ME6 Contextual as a baseline for our LLM-based classifiers.
@@ 3.1. EVP Online
The EVP Online11footnotemark: 1, developed by Cambridge University Press, provides CEFR levels for single words, phrasal verbs, phrases, and idioms. Each entry includes a PoS tag, a gloss, and both dictionary and learner examples. A key feature of the EVP is its sense-level CEFR annotation, which assigns a proficiency level to each sense. This level of granularity is beneficial for both general language education and the development of CALL systems. For this study, we used the single-word entries from the American English subset of the EVP, including their CEFR levels, PoS tags, glosses, and example sentences. Unlike a previous study Aleksandrova and Pouliot (2023) that also included multiword expressions (MWEs), our work focuses exclusively on single words, as WordNet contains very few MWEs.
@@ 3.2. SemCor Corpus
The SemCor 3.0 corpus is the most widely used sense-annotated corpus in NLP. It contains 226,040 sense annotations across 352 documents from the Brown Corpus. Each sense is tagged with a WordNet identifier, which links it to glosses, usage examples, and semantic relations such as hypernyms, hyponyms, synonyms, and antonyms. Moreover, its machine-readable format facilitates integration into NLP and CALL systems. However, this corpus inherits the limitations of WordNet mentioned in the introduction. Notably, the lack of learner-oriented indicators, such as sense complexity or CEFR levels, limits its utility for educational applications. To address this, as described in Section 5.1, we use our CEFR-annotated WordNet to enhance the original SemCor corpus, creating a new resource annotated with both senses and CEFR levels.
@@ 4. CEFR-Annotated WordNet
To create a WordNet oriented toward L2 learners, we annotated its senses with CEFR levels by aligning them with glosses from the EVP Online. The process, illustrated in Figure 2 for the WordNet gloss of word, PoSbank, noun comprises three steps:
@@
For each word and PoS pair, such as bank, noun, we extract all corresponding glosses from both the EVP Online and WordNet. Let the set of glosses from the EVP Online be , and that of glosses from WordNet be . In the next step, we focus on comparing the -th gloss, , from the EVP Online with the -th gloss, , from WordNet. As shown in Figure 2, both of the example glosses refer to sloping land.
@@
To measure the semantic similarity between and , we used an LLM (GPT-4o, checkpoint gpt-4o-2024-08-06). The prompt is shown in Appendix 13.1. Because glosses from different resources often vary in granularity and may not align perfectly, a binary alignment (same/different) judgment would be too restrictive. Therefore, we instructed the LLM to rate the similarity on a seven-point scale, where a lower score indicates higher similarity. In the example shown in Figure 2, the LLM returns a score of 1, indicating the two glosses have identical meaning.
@@
If the LLM returns a score of 1 or 2—indicating that and have “exactly the same” or “almost the same” meaning—we consider the glosses semantically aligned. The CEFR level associated with is then transferred to ; otherwise (i.e., output ), the senses are considered mismatched, and no annotation is made. In our example, the score of 1 results in the WordNet sense being assigned the B2 level from the corresponding EVP sense.
@@
whose size is . This exhaustive process yielded 10,995 CEFR-level annotations for 10,644 WordNet senses across 5,645 lemmas. Table 1 lists the distribution of these annotations. Nouns constitute the largest share (4,888; 44.46%), followed by verbs (3,163; 28.77%) and adjectives (2,327; 21.16%), with adverbs comparatively scarce (617; 5.61%). The distribution across CEFR levels is concentrated in the intermediate range, with B2 (29.30%) and B1 (20.77%) accounting for half of all annotations. The beginner (A1–A2) and advanced (C1–C2) levels represent 16.83% and 33.10%, respectively. Because the granularity of glosses differs between the two resources, a single WordNet sense can align with multiple EVP glosses, sometimes resulting in a sense being assigned multiple CEFR levels. This automated procedure is generalizable and could be applied to other dictionaries or lexical databases that provide glosses. However, because it is fully automated, evaluating the accuracy of the resulting annotations is essential.
5. Experiment
To verify the accuracy of our CEFR-level annotations, we built and evaluated several contextual lexical CEFR-level classifiers (Figure 1(b)). The goal was to assess how well models trained on our automatically annotated data could predict gold-standard CEFR levels. We also trained various LLM-based classifiers to evaluate their effectiveness for this task.

### v2

@@ Abstract
Although WordNet is a valuable resource because of its structured semantic networks and extensive vocabulary, its fine-grained sense distinctions can be challenging for second-language learners.
To address this issue, we developed a version of WordNet annotated with the Common European Framework of Reference for Languages (CEFR), integrating its semantic networks with language-proficiency levels.
@@
To validate our approach, we constructed a large-scale corpus containing both sense and CEFR-level information from the annotated WordNet and used it to develop contextual lexical classifiers.
Our experiments demonstrate that models fine-tuned on this corpus perform comparably to those fine-tuned on gold-standard annotations.
Furthermore, by combining this corpus with the gold-standard data, we developed a practical classifier that achieves a Macro-F1 score of 0.81. This result provides indirect evidence that the transferred labels are largely consistent with the gold-standard levels.
The annotated WordNet, corpus, and classifiers are publicly available to help bridge the gap between natural language processing and language education, thereby facilitating more effective and efficient language learning.
@@
| Masato Kikuchi✉††thanks: ✉ Corresponding author., Masatsugu Ono, Toshioki Soga , |
@@
| Nagoya Institute of Technology, Kitami Institute of Technology, |
| Chitose Institute of Science and Technology, Hokkaido University |
| {kikuchi, ozono}@nitech.ac.jp, onomasa@mail.kitami-it.ac.jp, |
@@ 1. Introduction
WordNet Fellbaum (1998) is a large-scale English lexical database that organizes approximately 155,000 words and 207,000 senses of nouns, verbs, adjectives, and adverbs into hierarchical semantic networks. It groups semantically similar words and links senses through relations such as hypernymy, hyponymy, synonymy, and antonymy. Because WordNet and its construction software are publicly available, they can be readily integrated into AI applications. Consequently, WordNet underpins a broad range of natural language processing (NLP) technologies—including machine translation Moussallem et al. (2018), semantic analysis Moskvoretskii et al. (2024), and natural language generation San Vicente et al. (2014)—owing to its accessible interface and well-structured networks. These technologies also support computer-assisted language learning (CALL) by facilitating vocabulary acquisition, reading comprehension, writing assistance, automated question generation, and automated assessment.
@@
Although leveraging semantic networks can enhance foreign-language learning Kiritani et al. (2012), WordNet was not designed for educational purposes and presents challenges for second-language (L2) learners. Key issues are its overly fine-grained sense distinctions and the large number of senses associated with many words. This requires learners to identify the appropriate sense for a given context and proficiency level, which increases their cognitive load. While this problem is widely discussed in NLP literature Navigli (2006)Lacerra et al. (2020), it has received limited attention in language education. Our goal was to develop a novel version of WordNet and leverage the resulting technologies and resources to enhance language-learning efficiency. The first step involves adapting WordNet for L2 learners and bridging the gap between NLP lexical resources and language education. Previous work Kikuchi et al. (2024) clustered fine-grained WordNet sense definitions (glosses) using learner-oriented dictionaries. By contrast, this study integrates Common European Framework of Reference for Languages (CEFR) proficiency levels into WordNet, enabling the presentation of senses aligned with a learner’s proficiency level. To build large-scale, practical resources, we employ a simple large language model (LLM)-based method for efficient and accurate semantic annotation, reducing the time, labor, and cost associated with manual annotation. This automatic approach also ensures that the adapted WordNet can be flexibly scaled.
@@
The CEFR is an international standard for describing language proficiency across six levels, namely, A1, A2, B1, B2, C1, and C2, ranging from basic to advanced. Each level is defined by “can-do descriptors” that specify expected communicative abilities. We used an LLM to annotate WordNet senses with CEFR levels, thereby constructing a CEFR-annotated WordNet. As shown in Figure 1(a), these annotations can be used in conjunction with semantic networks to help learners acquire vocabulary while considering relationships among words and learn basic and advanced paraphrases through synonyms. The annotation pipeline involves three steps. First, we collect glosses for target words from WordNet and the English Vocabulary Profile (EVP) Online111https://englishprofile.org/?menu=evp-online Capel (2012), which provides CEFR levels for individual senses. Next, the LLM computes the semantic similarity between the glosses from WordNet and the EVP. Finally, we assign CEFR levels to the corresponding WordNet senses based on these similarity scores.
@@
Because our method for annotating WordNet senses with CEFR levels is automatic, it eliminates the need for labor-intensive manual work. However, automatic labels may contain errors, and WordNet does not provide gold-standard CEFR levels. Therefore, their reliability must be verified indirectly. To address this issue, we built contextual CEFR-level classifiers that predict a sense’s proficiency level from its usage, as shown in Figure 1(b). These classifiers predict the CEFR level for a word sense based on its context, not just for the word itself. We evaluate the quality of our annotations by comparing classifiers trained on our data with those trained on the EVP gold-standard levels. We also examine the effectiveness of LLMs for this task in few-shot and fine-tuning settings.
@@
CEFR-Annotated WordNet. We developed a new resource by assigning CEFR proficiency levels to 10,644 WordNet senses corresponding to 5,645 lemmas, thereby linking the WordNet sense inventory with CEFR standards. Our annotated WordNet covers approximately 80% of all single-word senses in the EVP (8,289 out of 10,394).
Prompt-Only LLM Annotation. We introduce a novel, automated method that leverages the semantic understanding of LLMs to assign CEFR levels to word senses. This is achieved by measuring semantic similarity between WordNet glosses and EVP entries. The method, implemented entirely through prompting, is simple, reproducible, and substantially less costly than manual annotation. We also provide indirect evidence that manual annotation tasks based on semantic matching can be automated with high accuracy.
SemCor-CEFR Corpus. Using our annotated WordNet, we assigned CEFR levels to the word senses in SemCor 3.0222http://lcl.uniroma1.it/wsdeval/training-data Miller et al. (1993), a widely used sense-annotated corpus. This process resulted in a large-scale corpus containing more than 110,000 sense and level annotations across over 5,500 WordNet senses. Because modern NLP relies on large corpora for advanced training and analysis, this resource represents a valuable contribution to both NLP and educational technology research.
Contextual Lexical CEFR-Level Classifiers. We demonstrate the validity of our CEFR-level annotations by training a classifier on our corpus that performs comparably to one trained on gold-standard EVP data. Additionally, by fine-tuning the LLM on both our annotated data and the gold-standard levels, we developed a practical classifier that achieves a Macro-F1 score of 0.81. Our analysis indicates that these classifiers can accurately predict CEFR levels across a broad range of contexts.
@@ 2.1. WordNets for Language Learning
As noted in the introduction, WordNet was not originally designed for educational use. To address this limitation, several learning-oriented WordNets have been developed for multiple languages Bosch and Griesel (2018), and their application in language learning has been studied by many researchers Gonzalez-Dios (2019). Some of these researchers have focused on visualizing word hierarchies and semantic relations to support learners Sun et al. (2011); Kiritani et al. (2012); Gawde et al. (2024), while others have tailored vocabulary, glosses, and usage examples to match learners’ proficiency levels Redkar et al. (2018); Osenova and Simov (2024). However, most of these studies have relied on manually curated resources and paid limited attention to word-sense information. By contrast, herein we introduce a novel approach that automatically annotates WordNet senses with proficiency levels. Our method can be integrated with existing techniques—such as semantic network visualization and multimodal WordNets Marciniak (2020)—to further enhance its utility in language-learning contexts.
@@ 2.2. Lexical Complexity Prediction
Lexical complexity prediction (LCP) North et al. (2023); Shardlow et al. (2024) has recently attracted significant attention as a task for estimating word complexity from context. In this field, “complexity”—which is related to the CEFR levels in our study—is typically predicted as either binary (e.g., simple/complex) or on a continuous scale. Our work is closely related to SemEval-2021 Task 1 Shardlow et al. (2021), which adopts a similar classification setting. For this task, the organizers released the CompLex 2.0 dataset333https://github.com/MMU-TDMLab/CompLex Shardlow et al. (2022), in which words in context were rated by multiple annotators on a five-point Likert scale. The final scores are represented as a continuous value in , computed as the mean of these ratings. These continuous values can capture finer, context-driven differences compared with ordinal labels.
However, our approach differs from that of LCP in several key ways.
First, LCP is primarily designed as a precursor to lexical simplification Paetzold and Specia (2017)—which involves replacing complex words with simpler alternatives—rather than for explicitly presenting complexity information to L2 learners. Second, annotators of CompLex 2.0 were not provided with glosses; as a result, identical senses may receive different scores across contexts. Third, the dataset is limited to 9,000 nouns and excludes other parts of speech (PoS).
By contrast, our method assigns a CEFR level to each word sense in accordance with an international proficiency standard. We extend this annotation to more than 110,000 instances of nouns, verbs, adjectives, and adverbs in the large-scale SemCor corpus, making our resource more than ten times larger than CompLex 2.0. In Section 6.1, we analyze the correlation between the CompLex 2.0 complexity scores and the CEFR levels predicted by our models.
@@ 2.3. CEFR-Based Educational Technology
The CEFR is a foundational standard in educational technology and is widely applied in the automatic assessment of short sentences Tack et al. (2017); Uchida et al. (2024), teaching materials Pilán et al. (2016), writing skills Kerz et al. (2021); Schmalz and Brutti (2021), and learner proficiency Gaillat et al. (2022). The interaction between LLMs and the CEFR has been the focus of recent studies, which have explored how well these models understand proficiency levels Benedetto et al. (2025) and how to control the difficulty of the vocabulary and sentences they generate Alfter (2024); Malik et al. (2024); Barayan et al. (2025). These studies, together with the development of numerous CEFR-aligned lexical datasets, underscore the central role of the CEFR in the field.
For example, the CEFRLex project provides machine-readable lexical resources with word-level frequency counts by CEFR level Pintard and François (2020) for English and other languages444Only the Dutch resource NT2Lex provides frequency information at the word-sense level. Dürlich and François (2018); François et al. (2014); Tack et al. (2018); François et al. (2016); Volodina et al. (2016). However, it does not assign a unique CEFR level to each sense, making it unsuitable for tasks requiring sense-specific proficiency annotations. The Sense Complexity Dataset (SeCoDa) Strohmaier et al. (2020) provides sense-in-context CEFR annotations, but its sense labels are not aligned with those of WordNet. In addition, its small size, 1,432 words, limits its applicability within WordNet’s semantic framework. Our work addresses these gaps by annotating more than 110,000 word instances with sense-specific CEFR levels, thereby substantially expanding the available data on lexical difficulty.
Despite progress in LCP, few studies have examined the classification of vocabulary into CEFR levels based on context. Aleksandrova and Pouliot (2023) proposed ME6 Contextual, a BERT Devlin et al. (2019)-based classifier that, like our models, is trained on a CEFR-annotated corpus to directly predict a word’s level from its context. This direct prediction approach enables the model to classify words not seen during training. By contrast, Bannò et al. (2025) introduced an indirect-prediction method in which an LLM selects the appropriate EVP sense for a word in context and then maps it to a CEFR level. The performance of this approach depends on the quality of cues provided by the data source. To isolate the effect of different data sources on classification performance, we reimplemented ME6 Contextual as a baseline for our LLM-based classifiers.
@@ 3.1. EVP Online
The EVP Online11footnotemark: 1, developed by Cambridge University Press, provides CEFR levels for single words, phrasal verbs, phrases, and idioms. Each entry includes a PoS tag, a gloss, and both dictionary and learner examples. A key feature of the EVP is its sense-level CEFR annotation, which assigns a proficiency level to each sense. This level of granularity is beneficial for both general language education and the development of CALL systems. For this study, we used single-word entries from the American English subset of the EVP, including their CEFR levels, PoS tags, glosses, and example sentences. Unlike a previous study Aleksandrova and Pouliot (2023), which also included multiword expressions (MWEs), our work focuses exclusively on single words because WordNet contains very few MWEs.
@@ 3.2. SemCor Corpus
The SemCor 3.0 corpus is one of the most widely used sense-annotated corpora in NLP. It contains 226,040 sense annotations across 352 documents from the Brown Corpus. Each sense is tagged with a WordNet identifier, linking it to glosses, usage examples, and semantic relations such as hypernyms, hyponyms, synonyms, and antonyms. Moreover, its machine-readable format facilitates integration into NLP and CALL systems. However, this corpus inherits the limitations of WordNet discussed in the introduction. In particular, the absence of learner-oriented indicators, such as sense complexity or CEFR levels, limits its usefulness for educational applications. To address this issue, as described in Section 5.1, we use our CEFR-annotated WordNet to enhance the original SemCor corpus, creating a new resource annotated with both senses and CEFR levels.
@@ 4. CEFR-Annotated WordNet
To create a WordNet oriented toward L2 learners, we annotated its senses with CEFR levels by aligning them with glosses from the EVP Online. The process, illustrated in Figure 2 for the WordNet gloss of word, PoSbank, noun, comprises three steps:
@@
For each word and PoS pair, such as bank, noun, we extract all corresponding glosses from both the EVP Online and WordNet. Let the set of glosses from the EVP Online be , and that of glosses from WordNet be . In the next step, we compare the -th gloss, , from the EVP Online with the -th gloss, , from WordNet. As shown in Figure 2, both example glosses refer to sloping land.
@@
To measure the semantic similarity between and , we used an LLM (GPT-4o, checkpoint gpt-4o-2024-08-06). The prompt is provided in Appendix 13.1. Because glosses from different resources often vary in granularity and may not align perfectly, a binary alignment judgment of same or different would be overly restrictive. Therefore, we instructed the LLM to rate similarity on a seven-point scale, where a lower score indicates higher similarity. In the example shown in Figure 2, the LLM returns a score of 1, indicating that the two glosses have identical meanings.
@@
If the LLM returns a score of 1 or 2—indicating that and have “exactly the same” or “almost the same” meaning—we consider the glosses semantically aligned. The CEFR level associated with is then transferred to . Otherwise, that is, if the output is , the senses are considered mismatched and no annotation is assigned. In the example, the score of 1 results in the WordNet sense being assigned the B2 level from the corresponding EVP sense. In this study, we adopt a threshold of (scores 1–2) to balance accuracy and coverage. Restricting alignments to score 1 yields higher-confidence transfers but reduces coverage, whereas allowing score 3 increases coverage but tends to introduce false alignments, mainly because of partial semantic overlap and mismatched gloss granularity across resources. Therefore, we adopt as a conservative compromise.
@@
whose size is . This exhaustive process yielded 10,995 CEFR-level annotations for 10,644 WordNet senses across 5,645 lemmas. Table 1 lists the distribution of these annotations. Nouns constitute the largest share (4,888; 44.46%), followed by verbs (3,163; 28.77%) and adjectives (2,327; 21.16%), with adverbs comparatively scarce (617; 5.61%). The distribution across CEFR levels is concentrated in the intermediate range, with B2 (29.30%) and B1 (20.77%) together accounting for half of all annotations. The beginner levels (A1–A2) and advanced levels (C1–C2) represent 16.83% and 33.10%, respectively. Because the granularity of glosses differs between the two resources, a single WordNet sense may align with multiple EVP glosses, which can result in a sense being assigned multiple CEFR levels (see Appendix 13.2 for details). This automated procedure is generalizable and can be applied to other dictionaries or lexical databases that provide glosses. Moreover, because the pipeline relies solely on semantic similarity between annotations, it is transferable to lexical resources in other languages. However, as the process is fully automated, evaluating the accuracy of the resulting annotations is essential.
5. Experiments
To verify the accuracy of our CEFR-level annotations, we built and evaluated several contextual lexical CEFR-level classifiers (Figure 1(b)). The goal was to assess how well models trained on our automatically annotated data could predict gold-standard CEFR levels. We also trained several LLM-based classifiers to evaluate their effectiveness for this task. In this study, we did not conduct a manual spot-check of the transferred sense-level labels. Although expert validation would strengthen the reliability of the resource, rigorous CEFR-level judgment typically requires carefully designed protocols and multiple trained raters. We therefore leave such manual evaluation to future work.